whats the answer
solve each system of equations.
4y-3x=5
3/4x=y-4

Answer by fcabanski(1391)

Check my profile and contact me for one on one online tutoring.


When dealing with two equations and two variables, solve one equation for one of the variables, then substitute that into the other equation to solve for the variable. Then use that in either equation to solve for the second variable.


4y-3x=5
3/4x=y-4


Solve the first for y in terms of x: Add 3x to both sides:


4y=3x+5 Divide both sides by 4.


y = (3x+5)/4. Substitute that in for Y in the second equation:


3/4x =(3x+5)/4 - 4 Solve for X. First multiply all factors by 4.


3x=3x+5-16 Subtract x from both sides


0=-11 As written, there is no solution.